OpenStudy (anonymous):

the length of D rectangle is 5 inches more than the width.the perimeter is 70 inches. find the length. someone please help

OpenStudy (anonymous):

P = 2L + 2W L = W + 5 70 = 2(W + 5) + 2W 70 = 2W + 10 + 2W 70 = 4W + 10 60 = 4W 15 = W Now that we know the width, replace it in the original formula at the top of my post... 70 = 2L + 2(15) Solve. 70 = 2L + 30 40 = 2L 20 = L

OpenStudy (anonymous):

but it only says find the width does that mean i still have to do the second part and replace it

OpenStudy (anonymous):

Well in your question you said find the length. But if the question only asks for width, then you can stop at the 15 = W part.

OpenStudy (anonymous):

yes it actually did im sorry i was looking for another question. thank you so much for your help really appreciate it

OpenStudy (anonymous):

You're welcome (: